Sue can paint a barn by herself in X days. Mary can paint the same barn by herself in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together. Evaluate the expression for x = 5.
Answer by richwmiller(17219)

y/x+y/2x=1
2y+y=2x
3y=2x
y=2x/3
y/5+y/10=1
2y+y=10
3y=10
y=10/3
it takes them 10/3 (3 1/3 )days together if x=5
so they complete 3/10 in one day
so they complete 3x/2 in one day